OpenStudy (anonymous):

In a nuclear reaction, 2.50 x 1012 joules of energy have been released. What does this value equate to in kilograms? (1 J = 1 kg m2/s2) 8.33 x 104 kg 7.5 x 1020 kg 2.25 x 10-4 kg 2.78 x 10-5 kg

OpenStudy (anonymous):

Einstein's formula: \[E = mc ^{2}\] \[(2.50 x 10^{12} J) = (m)(3.00 x 10^{8} m/s)^2\] \[m = 2.78 x 10^{-5} kg\] In the future, could you show the exponent scientific notation like this: 2.50 x 10^12? (not 2.50 x 1012) It is a lot easier to read. Thanks.

OpenStudy (anonymous):

Yeah, sorry. Thank you.

OpenStudy (anonymous):

Where did the 3.00 come from? And the 10^8?

OpenStudy (anonymous):

Good question, that probably seemed liked it came out of nowhere. \[c\] is a constant that represents the speed of light. It is about 3.00 x 10^8 m/s. In the formula, it needs to be squared.

OpenStudy (anonymous):

Oh, okay thank you, again.

OpenStudy (anonymous):

Also what are the steps on how you got 'm' by itself?

OpenStudy (anonymous):

to start: (2.50x10^12 J) = (m)(3.00x10^8 m/s)^2 then square c: (2.50x10^12 J) = (m)(9.00x10^16 m/s) next divide both sides by c^2: (2.50x10^12 J)/(9.00x10^16 m/s) = (m) and the division yields the answer: m = 2.78x10^−5 kg (you know the units are in kg when they make that mention of 1 J = 1 kg m^2/s^2 at the end of the question.

OpenStudy (anonymous):

Okay, thank you, again.